Hi:
could anyone please see if they can solve "x":
-(x+7)/11 -(1-2x)/3=x
Thanks.

Answer by lwsshak3(11628)

solve for x
-(x+7)/11 -(1-2x)/3=x
LCD:33
-3(x+7)-11(1-2x)=33x
-3x-21-11+22x=33x
-14x=32
x=-32/14=-16/7